OpenStudy (lovelyharmonics):

Find the sum of the first 9 terms of the sequence. Show all work for full credit. 2, -7, -16, -25, ...

OpenStudy (anonymous):

It looks like the rule is: -9

OpenStudy (anonymous):

2-9=-7 -7-9=-16 -16-9=-25 -25-9=-34 -34-9=-43 -43-9=-52 -52-9=-61 -61-9=-70 So, the first 9 numbers in the sequence are 2,-7,-16,-25,-34,-43,-52,-61,-70

OpenStudy (anonymous):

So, just add all of those numbers to get your answer.

OpenStudy (anonymous):

2-7-16-25-34-43-52-61-70 = -306

OpenStudy (lovelyharmonics):

okay thank you c:

OpenStudy (anonymous):

Just combine the list I gave and caters answer to show full work. :)

OpenStudy (lovelyharmonics):

lol okay c: thankyou ^.^